# Rapid Rotation of Polarization Orientations in PSR B1919$+$21’s Single Pulses: Implications On Pulsar’s Magnetospheric Dynamics

### v1

@@ 5.1 Comments on modeling in Section 4
Some of our choices of parameters in Section 4.1 are mainly for simplicity. We let $\eta$ vary but $\delta$ keep constant—this assumption could be adjusted to let both $\eta$ and $\delta$ change, which could happen in real magnetospheres because of the difference on plasma particle distribution, for example. In addition, $E_{\mathrm{O}}/E_{\mathrm{E}}$, $\theta_{0}$ and $C$ need not to be constant along pulse longitudes. Our phenomelogical model could be applied to explain other polarization behaviours in pulsars, but we do not go into detail here.
@@
Eq. 1 is based on the assumption that the magnetic field is large enough. Quantitatively, validation of Eq. 1 requires $\omega_{B}\gg\omega$, where $\omega_{B}=eB/m_{e}c$. With dipole field $B\approx 10^{12}(R_{\mathrm{NS}}/r)^{3}\mathrm{G}$ we have $\omega_{B}=1.8\times 10^{19}(R_{\mathrm{NS}}/r)^{3}\mathrm{s}^{-1}$. So Eq. 1 is valid at least within $10^{3}R_{\mathrm{NS}}$ from the neutron star centroid. And because plasma particle number density decreases quickly when $r$ increases (Eq. 3), the integration of Eq. 2 will hardly differ between integrating to light cylinder and integrating to $\sim 10^{3}R_{\mathrm{NS}}$. The inner magnetosphere contributes to the phase lag most—this is also valid for oblique rotators ($\alpha>0$).
@@
$\eta$ in Section 4.2 is evaluated to be on the order of $2\pi$. To achieve rapid polarization rotations in single pulses, $\eta$ might be larger, which demands corresponding $\kappa$ to be larger or $\gamma$ to be smaller. But the requirement of low emission height should always be satisfied.
5.2 Bifurcated emission component in highly linearly polarized profile
In Figure 2, the highly linearly polarized profile has two emission components at the leading edge of the total integrated profile. The two components are close to each other, with a large depth in between. Such phenomenon is similar to the bifurcated emission component studied in Dyks et al. (2010); Dyks (2023). In some pulsars’ integrated profiles, bifurcated emission components are attributed to a curvature-radiation-induced split fan beam, where only the ordinary mode component is Doppler shifted into the observed bandwidth through the scattering (Dyks, 2023). We observe only one patch of OPM in Figure 2’s bifurcated component, which is also consistent with the model. The bifurcated component is invisible in the total integrated profile (Figure 2 (i)), where only a small precursor component is seen.
@@ 5.3 On asymmetries in B1919$+$21’s polarization patterns
We’d like to emphasize two asymmetries in our observational facts. One is that from the RVM fitting, the centroid of the PA curve seems to arrive manifestly earlier than the centroid (valley) of the integrated pulse’s profile of total intensity, which contradicts the model of relativistic abberation (BCW model, Blaskiewicz et al. 1991), assuming a symmetrically filled beam. The other, which is more rigid, is that most monotonic long PA curves (cover $>180^{\circ}$) in single pulses have a negative slope, while only very few of them have a positive slope. Both asymmetries could be interpreted as consequences of the asymmetric distribution of magnetospheric plasma particles’ number density or Lorentz factor, or of discharge positions near the pulsar surface. In the case of our theoretical explanation, PA slopes imply the increase or decrease of phase lag between O and E modes versus longitude, which depends on distributions mentioned above.
Considering the fact that PA curves could change greatly between adjacent pulses (see #274 and #275 in Figure 4 for examples), and that the dispersion relation (Eq. 1) depends sensitively on Lorentz factor $\gamma$, single pulses’ polarization pattern may mostly be affected by the $\gamma$ distribution of plasma particles in the magnetosphere. Distributions of particle number density and discharge positions seem more likely to be almost fixed, although they may be asymmetrical about the fiducial plane (where line of sight, magnetic axis and rotational axis are coplanar), probably due to some particular surface structures (e.g., Xu, 2024; Wang et al., 2024).
5.4 Why B1919$+$21 is “special”?
Although we have not achieved good RVM fitting in this paper, low emission height required in Section 4.2 seems to also support the alignment of magnetic axis and line-of-sight, which may also account for B1919$+$21’s relatively high brightness. The coherence of orthogonal wave modes, which could be related to the question where orthogonal modes are generated, needs further considerations.
The low $L/I$ in both B1919$+$21’s integrated profiles and single pulses also indicate significant incoherent superposition of waves. Such superposition may be related to the low emission height. To clarify the relation between emission height and wave superposition, we draw a schematic diagram of the radiative region above pulsar surface, in Figure 7 (left). We could receive radio waves from different emission points, where the radiation is concentrated within a $1/\gamma$ cone due to relativistic beaming. In the following discussion, we assume constant emission height $h$ and radiating particles’ Lorentz factor $\gamma$. When our line-of-sight is fixed, for example parallel to the magnetic axis, the emission points visible to us are determined by the positions and orientations of $1/\gamma$ cones. We mark the angle between line-of-sight and foot of magnetic field line, where the $1/\gamma$ cone’s inner edge is parallel to the line-of-sight, as $\phi_{e}$. The ratio of $\phi_{e}$ to the angular radius of polar cap $\phi_{p}$ can describe the degree of overlapping between different $1/\gamma$ cones. Figure 7 (right) shows $\phi_{e}/\phi_{p}$’s dependence on $h$ and $\gamma$. Smaller $h$ leads to larger $\phi_{e}/\phi_{p}$, which indicates more significant incoherent mixing of waves from different $1/\gamma$ cones.
Figure 7: Left: Schematic diagram for emission region above a pulsar’s polar cap. The grey and thin balck curves represent dipolar magnetic field lines. The thick circular arc means the pulsar surface. The orange sector means the radiation cone of $1/\gamma$. The orange arrow with label “$\boldsymbol{k}$” is the wave emission/propagation direction, which is tangential to local magnetic field. For other details, please refer to the text. Right: $\phi_{e}/\phi_{p}$’s dependence on $h/R_{\mathrm{NS}}$, plotted with 3 different values of $\gamma$ (100, 500 and 1000, marked with red, green and blue colors).
However, B1919$+$21 may not be “special” for its rapid rotating polarization orientations in single pulses, because many pulsars actually lack detailed single pulse investigations of polarization. A more refined model might be achieved after more similar or relevant phenomena are analyzed.
@@ 6 Conclusion
The FAST observational results of PSR B1919$+$21 are presented in this manuscript. The integrated profile’s complex PA curve and circular polarization variation result from the rapid rotating polarization orientations in single pulses. In many observed single pulses, the PA could monotonically rotate over $180^{\circ}$ or even $360^{\circ}$ across longitudes, and is always accompanied with circular polarization variations that could change handedness. Most of those quasi monotonic PA curves have negative slopes.
We attribute the phenomena to the difference in normal modes’ phase lags between different pulse longitudes, and build a simple model based on propagational effects that can reproduce the observed polarization patterns in single pulses. An appropriate value of phase lag between O and E waves requires appropriate values of pulsar magnetosphere’s multiplicity $\kappa$ and secondary plasma’s Lorentz factor $\gamma$. The emission height should be low enough to fit the weak frequency dependence of the polarization rotations. We argue that the distribution of plasma number density, of discharge points near pulsar surface, or of secondary particles’ Lorentz factor $\gamma$, should be asymmetric around the fiducial plane to explain the asymmetry observed in single pulses’ monotonic PA curves’ slopes.
Still we are not sure if B1919$+$21 is “special” for exhibiting such single pulse polarization pattern. Except for the requirement of more single pulse investigation of this pulsar under other frequency bands, we appeal that researchers do more single pulses studies on more known pulsars, for a better understanding of radio pulsars’ radiation and magnetospheres.
SSC wants to thank Prof. Gregory Herczeg at KIAA, Peking University. Through his “Advanced Writing for Astronomy” course, SSC’s writing of this paper had been improved. Shenglan Sun at Peking University also gave useful comments on the paper during the writing course by Gregory. JD acknowledges funding by the National Science Centre, Poland, grant no. 2023/49/B/ST9/01783. All data used in this work is from the Five-hundred-meter Aperture Spherical radio Telescope (FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work is supported by the National SKA Program of China (2020SKA0120100), the National Natural Science Foundation of China (Nos. 12003047 and 12133003), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0550300). Part of this research (Figure 1’s (ii)) has made use of the EPN Database of Pulsar Profiles maintained by the University of Manchester, available at: http://www.jodrellbank.manchester.ac.uk/research/pulsar/Resources/epn/.
@@ Appendix A Formula for constructing the phenomenological model
The mathematical details of the model described in Section 4.1 are presented in this section. We mark the initial O and E waves’ electric field vector amplitudes as $E_{O}$ and $E_{E}$. $E^{(C)}_{O,0}=C\cdot E_{O}$ and $E^{(C)}_{E,0}=C\cdot E_{E}$ represent the coherently added part, while $E^{(I1)}_{O,0}=(1-C)\cdot E_{O}$, $E^{(I1)}_{E,0}=0$ and $E^{(I2)}_{O,0}=0$, $E^{(I2)}_{E,0}=(1-C)\cdot E_{E}$ represent the incoherently added part (same as the settings in Oswald et al., 2023). After propagation in the first plasma layer, the electric field vectors become:
$E^{(C;I)}_{O,1}=E^{(C;I)}_{O,0}$ (A1)
$E^{(C;I)}_{E,1}=E^{(C;I)}_{E,0}e^{i\eta}$ (A2)
and after propagating through the second plasma layer, the electric field vectors become:
$E^{(C;I)}_{O,2}=E^{(C;I)}_{O,1}\cos{\theta_{0}}+E^{(C;I)}_{E,1}\sin{\theta_{0}}$ (A3)
$E^{(C;I)}_{E,2}=(-E^{(C;I)}_{O,1}\sin{\theta_{0}}+E^{(C;I)}_{E,1}\cos{\theta_{0}})e^{i\delta}$ (A4)
Then we return to the original coordinates, where the electric field vectors in the original O and E wave directions ($x$ and $y$ in Figure 5) are:
$E^{(C;I)}_{x}=E^{(C;I)}_{O,2}\cos{\theta_{0}}-E^{(C;I)}_{E,2}\sin{\theta_{0}}$ (A5)
$E^{(C;I)}_{y}=E^{(C;I)}_{O,2}\sin{\theta_{0}}+E^{(C;I)}_{E,2}\cos{\theta_{0}}$ (A6)
Finally we can calculate Stokes parameters (e.g., Jackson, 1998):
@@
The total Stokes parameters are:
@@
where $S=(I,Q,U,V)$. With Stokes parameters we can calculate PA and EA:
@@
$\mathrm{EA}=0.5\arcsin(V/\sqrt{Q^{2}+U^{2}+V^{2}})$ (A13)

### v2

@@ 5.1 Comments on modeling in Section 4
Some of our choices of parameters in Section 4.1 are mainly made for simplicity. We let $\eta$ vary but keep $\delta$ constant—this assumption could be adjusted to let both $\eta$ and $\delta$ change, which could happen in real magnetospheres because of the difference in plasma particle distribution, for example. In addition, $E_{\mathrm{O}}/E_{\mathrm{X}}$, $\theta_{0}$ and $C$ need not to be constant along pulse longitudes. In observed pulses, the $P/I$ curve seems to be related to the $I$ profile, which should be associated with coherence $C$. $C$ could also be different in two plasma layers. Some simulated pulses with different parameters, including one with varying $C$, are shown in Figure 8 in Appendix B. Our phenomenological model could be applied to explain other polarization behaviors in pulsars, but we do not go into detail here.
@@
Equation 8 is based on the assumption that the magnetic field is large enough. Quantitatively, validation of Equation 8 requires $\omega_{B}\gg\omega$, where $\omega_{B}=eB/m_{e}c$. With dipole field $B\approx 10^{12}(R_{\mathrm{NS}}/r)^{3}\mathrm{G}$ we have $\omega_{B}=1.8\times 10^{19}(R_{\mathrm{NS}}/r)^{3}\mathrm{s}^{-1}$. So, Equation 8 is valid at least within $10^{3}R_{\mathrm{NS}}$ from the neutron star centroid. And because plasma particle number density decreases quickly when $r$ increases (Equation 10), the integration of Equation 9 will hardly differ between integrating to the light cylinder and integrating to $\sim 10^{3}R_{\mathrm{NS}}$. The inner magnetosphere contributes most to the phase lag—this is also valid for oblique rotators ($\alpha>0$).
@@
$\eta$ in Section 4.2 is evaluated to be on the order of $2\pi$. To achieve rapid polarization rotations in single pulses, $\eta$ might be larger, which demands that the corresponding $\kappa$ to be larger or $\gamma$ to be smaller. But the requirement of low emission height should always be satisfied to achieve a small $\theta$ in Equation 8.
The phase-lag-driven model that we present performs well in understanding large rotations of PA, but the requirement of nearly equal mode strength ratio remains a question (see Figure 8 for a simulated pulse with $E_{\mathrm{O}}/E_{\mathrm{X}}=1.4$, where the monotonic rotation of PA ceases). We could also consider modeling that is based on the rotation of the emitted polarization orientation or on the variable mode amplitude ratio. The pure RVM-type PA curve can only cover $180^{\circ}$ of PA values. The change of wave modes amplitude ratio could also lead to rotations of polarization orientation, but if we keep the phase lag constant and consider only one pair of OPM, the PA variation will not be greater than $180^{\circ}$. After taking RVM into account, amplitude ratio-driven PA variation could cover $360^{\circ}$. This PA variation is still inadequate compared with some very large PA rotations (e.g. #2565 in Figure 4) of our observed single pulses. In order to achieve PA rotation over $360^{\circ}$, the emitted radio wave’s polarization orientation should initially be rotating at different longitudes before propagating, or there should be more pairs of OPMs added coherently. A separate paper is in preparation to discuss this.
The requirement of low emission height could be related to the possible small impact angle of viewing B1919$+$21. But B1919$+$21 may not be “special” for its rapid rotating polarization orientations in single pulses because many pulsars actually lack detailed single-pulse investigations of polarization. A more refined model might be achieved after more similar or relevant phenomena are analyzed.
@@ 5.2 On asymmetries in B1919$+$21’s polarization patterns
We would like to emphasize two asymmetries in our observational facts. One is that from the RVM fitting, the centroid of the PA curve seems to arrive manifestly earlier than the centroid (valley) of the integrated pulse’s profile of total intensity, which contradicts the model of relativistic aberration (Blaskiewicz-Cordes-Wasserman (BCW) model, Blaskiewicz et al. 1991), assuming a symmetrically filled beam. The other, which is more rigid, is that most monotonic long PA curves (cover $>180^{\circ}$) in single pulses have a negative slope, while only very few of them have a positive slope. Both asymmetries could be interpreted as consequences of the asymmetric distribution of magnetospheric plasma particles’ number density or Lorentz factor, or of discharge positions near the pulsar surface. In the case of our theoretical explanation, PA slopes imply the increase or decrease of phase lag between O and X modes versus longitude, which depends on the distributions mentioned above.
Considering the fact that PA curves could change greatly between adjacent pulses (see #274 and #275 in Figure 4 for examples), and that the dispersion relation (Equation 8) depends sensitively on the Lorentz factor $\gamma$, the polarization pattern of single pulses may mostly be affected by the $\gamma$ distribution of plasma particles in the magnetosphere. Distributions of particle number density and discharge positions seem more likely to be almost fixed, although they may be asymmetrical about the fiducial plane (where line of sight, magnetic axis, and rotational axis are coplanar), probably due to some particular surface structures (e.g., Xu, 2024; Wang et al., 2024).
@@ 6 Conclusion
The FAST observational results of PSR B1919$+$21 are presented in this manuscript. The complex PA curve and circular polarization variation of the integrated profile result from the rapid rotations of polarization orientations in single pulses. In many observed single pulses, the PA could monotonically rotate over $180^{\circ}$ or even $360^{\circ}$ across longitudes, and is always accompanied by circular polarization variations that could change handedness. Most of those quasi-monotonic PA curves have negative slopes.
We attribute the phenomena to the difference in normal modes’ phase lags between different pulse longitudes, and build a simple model based on propagational effects that can reproduce the observed polarization patterns in single pulses. An appropriate value of phase lag between O and E waves requires appropriate values of pulsar magnetosphere’s multiplicity $\kappa$ and secondary plasma’s Lorentz factor $\gamma$, which could be related to magnetosphere compositions. The emission height should be low enough to fit the weak frequency dependence of the polarization rotations. We argue that the distribution of plasma number density, of discharge points near the pulsar surface, or of secondary particles’ Lorentz factor $\gamma$, should be asymmetric around the fiducial plane to explain the asymmetry observed in single pulses’ monotonic PA curves’ slopes. Further explorations into particle acceleration processes and pulsar surface properties may help further our understanding of them.
Still, we are not sure if B1919$+$21 is “special” for exhibiting such a single-pulse polarization pattern. Except for the requirement of more single-pulse investigation of this pulsar under other frequency bands, we appeal that researchers do more single-pulses studies on more known pulsars for a better understanding of radio pulsars’ radiation and magnetospheres.
S.S.C. thanks Prof. Gregory Herczeg at KIAA, Peking University. Through his “Advanced Writing for Astronomy” course, S.S.C.’s writing of this paper was improved. Shenglan Sun at Peking University also gave useful comments on the paper during the writing course by Prof. Gregory Herczeg. An anonymous reviewer is acknowledged for providing us with many suggestions. J.D. acknowledges funding by the National Science Centre, Poland, grant no. 2023/49/B/ST9/01783. All data used in this work is from the FAST (Five-hundred-meter Aperture Spherical radio Telescope) (https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work is supported by the National SKA Program of China (2020SKA0120100), the National Natural Science Foundation of China (Nos. 12003047 and 12133003), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0550300).
@@ Appendix A Formula for constructing the phenomenological model
The mathematical details of the model described in Section 4.1 are presented in this section. We mark the initial O and X waves’ electric field vector amplitudes as $E_{\mathrm{O}}$ and $E_{\mathrm{X}}$. $E^{(C)}_{\mathrm{O},0}=C\cdot E_{\mathrm{O}}$ and $E^{(C)}_{\mathrm{X},0}=C\cdot E_{\mathrm{X}}$ represent the coherently added part, while $E^{(I1)}_{\mathrm{O},0}=(1-C)\cdot E_{\mathrm{O}}$, $E^{(I1)}_{\mathrm{X},0}=0$ and $E^{(I2)}_{\mathrm{O},0}=0$, $E^{(I2)}_{\mathrm{X},0}=(1-C)\cdot E_{\mathrm{X}}$ represent the incoherently added part (same as the settings in Oswald et al., 2023). After propagation in the first plasma layer, the electric field vectors become
$E^{(C;I)}_{\mathrm{O},1}=E^{(C;I)}_{\mathrm{O},0}$ (A1)
$E^{(C;I)}_{\mathrm{X},1}=E^{(C;I)}_{\mathrm{X},0}e^{i\eta}$ (A2)
and after propagating through the second plasma layer, the electric field vectors become
$E^{(C;I)}_{\mathrm{O},2}=E^{(C;I)}_{\mathrm{O},1}\cos{\theta_{0}}+E^{(C;I)}_{\mathrm{X},1}\sin{\theta_{0}}$ (A3)
$E^{(C;I)}_{\mathrm{X},2}=(-E^{(C;I)}_{\mathrm{O},1}\sin{\theta_{0}}+E^{(C;I)}_{\mathrm{X},1}\cos{\theta_{0}})e^{i\delta}$ (A4)
Then, we return to the original coordinates, where the electric field vectors in the original O and X wave directions ($x$ and $y$ in Figure 6) are
$E^{(C;I)}_{x}=E^{(C;I)}_{\mathrm{O},2}\cos{\theta_{0}}-E^{(C;I)}_{\mathrm{X},2}\sin{\theta_{0}}$ (A5)
$E^{(C;I)}_{y}=E^{(C;I)}_{\mathrm{O},2}\sin{\theta_{0}}+E^{(C;I)}_{\mathrm{X},2}\cos{\theta_{0}}$ (A6)
Finally, we can calculate Stokes parameters (e.g., Jackson, 1998):
@@
The total Stokes parameters are
@@
where $S=(I,Q,U,V)$. With Stokes parameters, we can calculate PA and EA
@@
$\mathrm{EA}=0.5\arctan(V/\sqrt{Q^{2}+U^{2}})$ (A13)
@@
Appendix B Six simulated pulses
Six simulated pulses with different groups of parameters are shown in Figure 8.
Figure 8: Six pulses simulated from the phenomenological model. The meanings of the dots and lines are the same as those in Figure 6 (iv). The parameters used to simulate the pulses are given below each plot.